OpenStudy (anddrea_x33):

Will Fan and Medal! please help. The Sugar Sweet Company is going to transport its sugar to market. It will cost $5500 to rent trucks, and it will cost an additional $225 for each ton of sugar transported. Let C represent the total cost (in dollars), and let S represent the amount of sugar (in tons) transported. Write an equation relating C to S. Then use this equation to find the total cost to transport 19 tons of sugar.

OpenStudy (anddrea_x33):

equation: total cost to transport 19 tons of sugar: $

OpenStudy (jonaho):

C + (S x 175) = t (t stands for the total) 4500 + (19 x 175) = t 4500 + 3325 = t $7825 = t

OpenStudy (kendricklamar2014):

@anddrea_x33 next time put this question in the Math Section. Since this is the OS Feedback section, you will get even more help in the Mathematics section. \(\Large Good~Luck~and~Welcome~to~OpenStudy\)

OpenStudy (confluxepic):

Welcome to OpenStudy! @anddrea_x33